The vapor pressure of pure water at 85 °C is 434 torr. What is the vapor pressure at 85 °C of a solution prepared from 100 mL of water (density 1.00 g/mL) and 150 g of diglyme, C6H14O3, a nonvolatile substance?


1
Expert's answer
2021-04-28T02:59:39-0400

Note: Diglyme is the solute and water is the solvent.

Calculating the number of moles of each component

n solute = 150 g × (1 mol / 134.17 g) = 1.1180 mol

n solvent = 100 g × (1 mol / 18.015 g) = 5.5509 mol

Calculating the total number of moles

n T = n solute + n solvent

n T = 1.1180 mol + 5.5509 mol

n T = 6.6689 mol

Calculating the mole fraction of the solvent

X solvent = n solvent / n T

X solvent = 5.5509 mol / 6.6689 mol

X solvent = 0.83236

Calculating the vapor pressure of the solution

P solution = (X solvent)(P solvent)

P solution = (0.83236)(434 mmHg)

P solution = 361 mmHg